Present: Eric Myers, Tibi, Liz and Nick

Ligo Analysis and the Grid

Eric and Tibi had a long discussion of the requirements for the LIGO analysis and related grid issues. They will follow up via phone later this week.

Cosmic Analysis

Nick uncovered issue with the error checking part of the analysis beans is not being rolled out. Ben changed the way to do it, but it got lost somehow. Ben is going to find the code and fix it.

Cosmic Analysis

Load-balancing - 14,15,16,17 ready; also can use 10 Tibi will test with Jmeter.

Problems with Cosmic e-Lab

Tom Jordan had to restart Tomcat twice during a workshop this week. He noticed that data1 was getting an extremely heavy load. Why does it get bogged down? Mihael wants a new hard drive for data1. Tibi wants to clean up data0. Mihael thinks the RLS is buggy. Tibi thinks that RLS may run into problems because data0 is so full. Load balancing may be big help for workshops. Bob is concerned that the NSF mounts may be slow; Action Item: Tibi will communicate with Ben and Mihael about cleaning up data0


We have improvements in analysis and minor fix for teacher registration to allow alphanumeric data for schools.

Rollout ? Tibi will work with Nick on this.

-- Main.LizQuigg - 13 Jun 2007
-- Main.EricMyers - 14 Jun 2007
Topic revision: r4 - 2007-06-14, EricMyers
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback